if x=a+b,y=a ω +b ω 2 ,z=a ω 2 +b ω , then find the value of xyz ?


if x=a+b,y=aω+bω2,z=aω2+bω, then find the value of  xyz ?


Grade:

1 Answers

Kevin Nash
14 years ago

Dear Mahima,

As you know that 'w' is the cube root of unity:

 so w3 = 1           &            w2+w + 1 = 0;

so 

         (a+b)(aw+bw2)(aw2+bw)

       = w2 (a+b)(a+bw)(aw+b)

        =(a+b)(a+bw)(aw3+bw2)

        =(a+b)(a+bw)(a+bw2) =a3+b3

  Thus your required answer is  : " a3+b3 "

askIITians.com provides online iit jee courses and IIT JEE Test Series with IITians. Click here to get free online test series and check your status timely or you can join us as our registered user for getting best iit jee study material or iit jee test series.

Click here to find IIT JEE Packages or course details with askIITians.

for further details visit the folowing links:

http://www.askiitians.com//iit-jee-online-coaching.aspx

http://www.askiitians.com/iit-jee-test-series.aspx

http://www.askiitians.com/iit-study-material/

Askiitians provide you all the best possible information, feel free to ask.

all the best.

thanks and regards.

Akhilesh Shukla